OpenStudy (anonymous):

7x + 3y = 30 -2x + 3y = 3 A. (3, 5) B. (6, 3) C. (3, 3) D. (6, 5)

OpenStudy (anonymous):

7x + 3y = 30 ----> (-1) -2x + 3y = 3 -------------- -7x - 3y = -30 (result of multiplying by -1) -2x + 3y = 3 --------------add -9x = - 27 (divide both sides by -9) x = 3 now sub 3 in for x in either equation to find y 7x + 3y = 30 7(3) + 3y = 30 21 + 3y = 30 3y = 30 - 21 3y = 9 y = 3 check.. -2x + 3y = 3 -2(3) + 3(3) = 3 -6 + 9 = 3 3 = 3 (correct) x = 3 and y = 3 or (3,3)